A certain group of test subjects has pulse rates with a mean of 84.2 beats per minutes and a standard deviation of 10.8 beats per minutes. Would it be unusual for one of the test subjects to have a pulse rate of 95.8 beats per minute?
2. (15) A certain group of test subjects has pulse rates witha mean of 84.2 beats per minutes and a standard deviation of 10.8 beats per minutes. Would it be unusual for one of the test subjects to have a pulse rate of 95.8 beats per minute? A. Minimum usual value beats per minute B. Maximum usual value beats per minute C. Is 95.8 beats per minute an unusual pulse rate? (Circle the correct response.) No, because it is larger than the maximum usual value. Yes, because it is larger than the maximum usual value. Yes, because it is between the minimum and maximum usual values. Yes, because it is smaller than the minimum usal value. No, because it is between the minimum and maximum usual values. No, because it is smaller than the minimum usual value
Answer #1

Q) Given that,

Mean (mu) = 84.2 beats per minutes

Standard deviation (sigma) = 10.8 beats per minute

μ-20-84.2-(2 * 10.8) = 62.6

μ + 2σ = 84.2 + (2 * 10.8) 105.8

A) Minumum "usual" value = 62.6 beats per minute

B) Maximum "usual" value = 105.8 beats per minute

C) Here, 95.8 beats per minute is not an unusual pulse rate. Because, 62.6 < 95.8 < 105.8

Answer: No, because it is between the minimum and maximum "usual" values.